Home‎ > ‎Crew 131 - Embry Riddle‎ > ‎

1211-EngineeringReport

Engineering Report
12/11/2013 Evening

Chui Lee
Was Power Interrupted:
no
Time Of Power Outage:

Length Of Power Outage:

Diesel Fuel Level:
.5
Propane Fuel Level:
.27
Gasoline Fuel Level:
1.0
Fuel Notes And Comments:

Atv Oil Quarts:
0
Car Oil Quarts:
3.5
Oil Notes And Comments:

Trailer Potable Water Level:
26.035
Trailer Potable Water Refilled:
no
Outside Potable Water Level:
28.575
Outside Potable Water Refilled:
no
Loft Potable Water Gallons:
35
Loft Potable Water Refilled:
no
Potable Water Meter Reading:
70702.1
Greenhouse Notes And Comments:
Heater/pilot light fully functioning, GreenHab at 62 Degrees F
Telescope Used:
no
Observatory Notes And Comments:

Network Failed:
no
Network Throttled:
no
Network Signal Strength:
194 (14.24dB)
Network Download Allowane:
16
Network Notes And Comments:

Hvac Air Conditioner Used:
no
Hvac Furnace Used:
yes
Hvac Lab Wall Heater Used:
no
Hvac Upper Deck Thermostat Setting:
75
Hvac Upper Deck Temperature Reading:
71
Hvac Lower Deck Thermostat Setting:
75
Hvac Lower Deck Temperature Reading:
52.5
Hvac Notes And Comments:

Hab Car Used:
no
Hab Car Oil Checked:
no
Hab Car Tire Status:
Good
Hab Car Odometer:
130658.3
Hab Car Notes And Comments:

Atv One Used:
no
Atv One Oil Checked:
no
Atv One Fuel Used Gals:
.5
Atv One Tires Status:
Good
Atv One Hours Used Day:
.75
Atv One Hours Used Year:
.75
Atv One Notes And Comments:

Atv Two Used:
no
Atv Two Oil Checked:
no
Atv Two Fuel Used Gals:
.5
Atv Two Tires Status:
Good
Atv Two Hours Used Day:
.75
Atv Two Hours Used Year:
1.58
Atv Two Notes And Comments:
ATV was rolled over. Conducted assessment and there are no cosmetic nor mechanical flaws due to the accident.
Atv Three Used:
no
Atv Three Oil Checked:
no
Atv Three Fuel Used Gals:
.5
Atv Three Tires Status:
Good
Atv Three Hours Used Day:
.75
Atv Three Hours Used Year:
1.58
Atv Three Notes And Comments:
ATV swerved, but suffered no damage.
Atv Four Used:
no
Atv Four Oil Checked:
no
Atv Four Fuel Used Gals:
0
Atv Four Tires Status:
Good
Atv Four Hours Used Day:
0
Atv Four Hours Used Year:
.75
Atv Four Notes And Comments:
Battery is dead. Will bring in to charge tonight.
Atv Five Used:
no
Atv Five Oil Checked:
no
Atv Five Fuel Used Gals:

Atv Five Tires Status:

Atv Five Hours Used Day:

Atv Five Hours Used Year:

Atv Five Notes And Comments:

Comments:
(Notes sent in earlier email):
EVA Suits/Helmet:
Today during one of the EVAs, one of the crew members spun out, causing another crew member to swerve and flip their ATV. Because the EVA helmets were attached to the back of the ATVs, one of the helmets slipped out while one got crushed under the TV that flipped. As a result the bulbs (front face) of the helmets were destroyed, along with the frame of the helmets themselves (see attached). The crew members only brought back the backing of the helmets, as the front faces were not salvageable.

>> Please note-- Both crew members are in good condition. One suffered some bruising/scratches on his elbow, and soreness in his calf. I have wrapped up his arm and will be keeping an eye on his leg. For the time being we will be limiting his workload.

>> I am seeking permission to try and repair or pair up some of the air tubes that were attached to the helmets to working packs, as some of the packs are currently running weak (to which I am still investigating a fix). Because the tubes were mis-matched before our crew came, I will likely be trying to match the correct tubes to their respective packs. This has been an issue since we arrived, but I have been occupied with the internet issue, pipes and other packs to change the tubes. In terms of the broken helmets, I am unable to find parts to replace the front face-- what does Mission Control recommend as further action?

>> Water Pump/Pipes:
DG came and assessed the tube. He agreed that the seal/adapter needed to be replaced. He returned home to retrieve the part and is likely to return in approximately 2 hours. Currently the pump is detached from the main HAB pipes, thus there is no leak. DG also brought the trailer tank and filled the static tank to full.


New Notes:
>> Please note-- Both crew members are in good condition. One suffered some bruising/scratches on his elbow, and soreness in his calf. I have wrapped up his arm and will be keeping an eye on his leg. For the time being we will be limiting his workload.

>> In order to fix the weak packs, this may require replacement filters. We are unable to find filters around the HAB. I have look at the wiring of the packs and they all seem as solid as they can be, concerning the extent of their usage, so I do not believe this is the cause of the issue. Requesting feedback for packs or replacement filters.